The unit digit of a two-digit number is 3 and seven times the sum of the digits is the number itself. Find the number.

Answer by Edwin McCravy(20059)

The unit digit of a two-digit number is 3
u = 3
and seven times the sum of the digits is the number itself.
7(t+u) = 10t+u

Substitute 3 for u:

7(t+3) = 10t+3
 7t+21 = 10t+3
   -3t = -18
     t = 6
Find the number.
The tens digit is 6 and the units digit is 3, so the number is 63.

Edwin